In 2002, Toronto General Hospital telephone interviewed 1500 people and 645 support stem cell research. The same study was carried out in 2009 and 870 out of 1500 respondents support stem cell research. Is there any evidence to suggest a change on people’s view on stem cell research has changed?
Answer #1

Here we have given that,

Claim: To check whether the change on peoples view on stem cell research has changed or not.

The Hypotheses is

Ho: P1=P2

v/s

H1:P1 \neq P2

we have given that,

n1= number of interviewed people in 2002=1500

x1=number of people support stem cell research=645

n2=number of people interviewed in 2009=1500

X2=number of people support stem cell research =870

Now we estimate the proportion p as

\hat p 1=1st sample proportion = \frac{x1}{n1}=\frac{645}{1500} =0.43

\hat p 2= 2nd sample proportion =\frac{x2}{n2}=\frac{870}{1500} =0.58

Now, we can find the test statistics

Zstat=\frac{\hat p1 - \hat p2 }{\sqrt { \frac{\hat p1 *(1-\hat p1))}{n1} +\frac{\hat p2 *(1-\hat p2)}{n2}} }

=\frac{0.43 - 0.58 }{\sqrt {\frac{0.43 *(1-0.43))}{1500}+\frac{0.58 *(1-0.58)}{1500}}}

= - 8.31

Now

Pvalue = 0.0001 ( using excel =normsdist(zstat))

Here Pvalue< 0.05 \alpha

Conclusion:

That is here there is sufficient evidence that the change on peoples view on stem cell research has changed